OpenStudy (anonymous):

Would someone that knows how to simplify exponents tell me if my homework answers are correct? Here is the first one and my answer: 16^-5/4 my answer 1/32 Second question: (1/4) ^3/2 my answer: 1/128 Am I correct on both questions? If not please lead me to the answers. Thanks!

OpenStudy (anonymous):

Help! How do I figure the second one?

OpenStudy (anonymous):

The first answer is correct. The second answer is incorrect. To raise a number to the 3/2 power, we take the square root of it, and then cube that answer (it actually doesn't matter what order you do these in). So, \[.25^{3/2} = (\sqrt{.25})^3\]The square root of 1/4 is 1/2, so \[(\sqrt{.25})^{3} = (1/2)^{3} = 1/(2^{3}) = 1/8\] So our answer is 1/8.

OpenStudy (anonymous):

Thank you Vinny, I am copying your strategy for my notes. Can you tell me if this answer is correct for this problem? √50xz^10 my answer 3. ^5√(2 √xz^10

OpenStudy (anonymous):

The three is not part of the answer.

OpenStudy (anonymous):

The way you typed the question and answer there is a little unclear, could you try using the equation button below the box you're typing in to set it up a little more clearly?

OpenStudy (anonymous):

Here is the question: √(50xz^10)

OpenStudy (anonymous):

my answer is

OpenStudy (anonymous):

^5√(2 √(xz^10

OpenStudy (phi):

if that a square root or 5th root ?

OpenStudy (anonymous):

ok, so the question is \[\sqrt{50x(z^{10})}\] You can pull out a 5 in front, which I see you did. So now you have \[\sqrt{2x(z^{10})}\] But theres one more step you can do. You have z^10 which you can take the square root of. When you square root something, you are halving the exponent, so the square root of z^10 is z^5, which you can pull in front, so your final answer is \[5(z^{5})\sqrt{2x}\]

OpenStudy (anonymous):

I understand and will copy your notes! You deserve a medal, here you go!

OpenStudy (anonymous):

Thank you, I'm glad I could help.

OpenStudy (anonymous):

You have definitely helped, I get really nervous with homework assignments!

OpenStudy (anonymous):

I have a new one can you help me with this one? I have never seen this in class before. (5√3-2√6) (2√6+√3)

OpenStudy (anonymous):

The answer in the calculator is 24√2-9

OpenStudy (anonymous):

The equation states to multiply.

OpenStudy (anonymous):

Ok. We need to mutlply these two together, and we will use the FOIL method. Are you familiar with this method? It's a rather simple way to multiply two pieces of the form (a+b)(c+d). When we foil this out we get - First - (5√3)(2√6) = 10√18 = 30√2 Outside - (5√3)(√3) = 5√9 = 15 Inside - (-2√6)(2√6) = -4√36 = -24 Last - (-2√6)(√3) = -2√18 = -6√2 So our final answer is 30√2 + 15 - 24 - 6√2

OpenStudy (anonymous):

We can simplify this to 24√2 - 9.

OpenStudy (anonymous):

It does ask to simplify as much as possible! Thank your for writing the steps!

OpenStudy (anonymous):

Are you up for a harder equation it say to simplify?

OpenStudy (anonymous):

a^5/2 a^-1/2/a^1/3 the a^1/3 is the denominator to rest is the numerator.

OpenStudy (anonymous):

Ok, so it looks like this? \[\frac{ a^{5/2} \times a^{-1/2} }{ a^{1/3} }\]

OpenStudy (anonymous):

The denominator is a ^1/3 Now there was not any implication to multiply no x or dot to imply multiplication, it states to simplify. I got a^25/6/2

OpenStudy (anonymous):

If they're next to each other in the numerator then they are multiplied. When we multiply, we add the exponents, when we divide, we subtract them. \[\frac{ a^{(5/2)-(1/2)} }{ a^{1/3} } = \frac{ a^{4/2} }{ a^{1/3} } = a^{4/2 - 1/3} = a^{5/3}\]

OpenStudy (anonymous):

Clearly I am algebra challenged now does it matter that the negative symbol is next to -1/2 it is not outside? I mean would that change the answer? Thank you for your help!

OpenStudy (anonymous):

Shouldn't it be -(-1/2) and would that change the answer?

OpenStudy (anonymous):

-(1/2) and -1/2 are equivalent. In the first fraction I put, you're actually adding 5/2 + (-1/2) but I skipped a step and wrote it as 5/2 - 1/2. Sorry if I confused you.

OpenStudy (anonymous):

Hi Vinny, I was all ready confused. In the example the denominator is multiplied with the denominator of the numerator, wouldn't that be 3 x 2 = 6?

OpenStudy (anonymous):

The example only has one fraction at the top so it is not completely similar and is throwing me off!

OpenStudy (anonymous):

The base is always the same, so when we multiply together, we add the exponents. The numerator had some multiplication so thats why we add the exponents. After I simplified the numerator, we need to evaluate a^(4/2) / a^(1/3). Since we are dividing, we must subtract the exponents. 4/2 - 1/3 is equal to 5/3, which becomes our exponent.

OpenStudy (anonymous):

So the final answer is a^5/2 correct? P.S. How would I write b^4/7 as a radical expression?

OpenStudy (anonymous):

The final answer would be a^5/3. This is assuming I understood the question you posed to me correctly. When we have a fractional exponent, the number of the exponent is the power, and the denominator is the root. So, for b^(4/7) we have \[b^{4/7} = (\sqrt[7]{b})^{4}\]

OpenStudy (anonymous):

Thank you for that because I wrote it backwards. How do you process the steps for this?

OpenStudy (anonymous):

Do I need the parentheses? and is this correct: (7√b )^4

OpenStudy (anonymous):

Thank you for the explanation. I have asked this question in an e-mail to my teacher and never got a response. Thank you for that!

OpenStudy (anonymous):

The parenthesis are actually not necessary because the following are all equivalent: \[b^{4/7} = (\sqrt[7]{b})^{4} = \sqrt[7]{(b^{4})}\]

OpenStudy (anonymous):

Last question how would I simply this? √50/32 The radical symbol stretches the length of the fraction.

OpenStudy (anonymous):

The calculator shows 5/16√2

OpenStudy (anonymous):

or 5/4

OpenStudy (anonymous):

If the radical stretches over the whole fraction, then the answer can be found by finding the radical of the top over the radical of the bottom. \[\sqrt{50/32} = \sqrt{50} / \sqrt{32} = \frac{ 5\sqrt{2} }{ 4\sqrt{2} } = \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

Thank you for showing your work! I am finding that I am actually learning in this format! Thank you for being such a good teacher!

OpenStudy (anonymous):

Please check my first equation: 8w√(48wu^2 –u √75^3) my answer: 27uw√3w

OpenStudy (anonymous):

Vinny I hope that you do not mind checking homework with me. I forgot to ask if you had the time!

OpenStudy (anonymous):

Ok, does it look like this? \[(8w)\sqrt{48wu^{2} - (u)\sqrt{75^3}}\]

OpenStudy (anonymous):

There are two radical signs: 8w√48wu^2 -u √75^3

OpenStudy (anonymous):

it is 75 cubed

OpenStudy (anonymous):

75w^3

OpenStudy (anonymous):

Hows this - \[8w \sqrt{48wu^2} - \sqrt{75w^3}\]

OpenStudy (anonymous):

Hows this - \[8w \sqrt{48wu^2} - \sqrt{75w^3}\]

OpenStudy (anonymous):

There is a -u in the center but all everything else is perfect!

OpenStudy (anonymous):

Vinny is 27uw √3w correct?

OpenStudy (anonymous):

Ok I see what you mean. Sorry for the delay, I'm tackling way too many questions at once. \[8w \sqrt{48wu^2} - u\sqrt{75w^3}\] We can pull out a 16 from the 48, which will give us a 4 in front and leave a 3 in the radical. We can also square root the u^2. From the second radical pull out and square root the 25. \[32wu \sqrt{3w} - 5u\sqrt{3w^3}\] Now, we can pull out a w^2 and square root it. Simplify and we're done. \[32wu \sqrt{3w} - 5wu\sqrt{3w} = 27uw \sqrt{3w}\] Looks good!

OpenStudy (anonymous):

Yeah! My last question, and I just want to know if my answer is correct. 2. Rationalize the denominator and simplify: √5/√13 answer √65/13

OpenStudy (anonymous):

Multiply top and bottom by √13, and you're done. You've got it.

OpenStudy (anonymous):

Ok, I multiplied 5 x 13 and got 65 squared over 13, am I correct?

OpenStudy (anonymous):

Yep, thats perfect. You definitely have a good handle on this stuff.

OpenStudy (anonymous):

Awwwww! You make me feel so much better because I am seriously a novice!

OpenStudy (anonymous):

With a little practice you can quickly become a pro. You could look online for some more practice problems if you'd like to build your confidence. I think I'm signing off for the night though.

OpenStudy (anonymous):

One more check if you do not mind? If not THANK YOU! Checking this question: Solve :(v-1)^3-88=0 answer v=1+2^3√11

OpenStudy (anonymous):

THANKS AGAIN VINNY FOR ALL OF YOUR HELP!

OpenStudy (anonymous):

The answer that you should be getting is \[\sqrt[3]{88} +1\] By adding 88 to both sides, cube rooting both sides, and adding 1 to both sides. I'm signing off for the night. If I've been helpful and you think I deserve it, I would greatly appreciate you adding me as a fan or even writing a testimonial on my profile. I'm glad to help.

OpenStudy (anonymous):

What is the number outside of the radical symbol?

OpenStudy (anonymous):

Thats a 3, it's the cube root of 88. You could pull out an 8 and cube root it. I think thats what you meant to say it was just typed a little inaccurately.

OpenStudy (anonymous):

Ok I am a fan and you have been great. When I finish my homework I will write a testimonial because you not only guided me, you helped to teach me concepts I had a hard time learning in the classroom. That's a good testimony! HAVE A GREAT EVENING!

OpenStudy (anonymous):

Thanks! Have a great weekend.

OpenStudy (anonymous):

I WILL NOW THANKS TO YOU! Hope to work with you again! Tuiti